OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, -7) and (6, -5)?

jimthompson5910 (jim_thompson5910):

Similar Example Find the slope of the line through (-3, -2) and (8, -4) m = (y2-y1)/(x2-x1) m = (-4-(-2))/(8-(-3)) m = (-4+2)/(8+3) m = -2/11 So the slope of the line that passes through the two points (-3, -2) and (8, -4) is -2/11 Note: this is NOT the answer as it's just an example

OpenStudy (anonymous):

Can i ge the answer im kind of in ahurry